OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −5 −6 −14 14 6

OpenStudy (anonymous):

All you need to do is plug in your terms. 4(-1)-2(-5) = -4+10 = your answer.